Assuming complete dissociation of H2SO4 as: (H_2SO_4+2H_2O \rightarrow 2H_3O^+SO_4^{2-})The number of H3O+ ions in 10 ml of 0.5 N H2SO4 is

        a) 3.01 × 1021 c) 6.02 × 1021

 

Grade:

3 Answers

Vishnu
4 years ago
Imole=6.022*10^23
In one mole of H3o 0.5moles are there
So 
 
 
6.022*10^23*0.5/2=3.01*10^21/2
~1.5*10^23
 
😉
 
And, :A
Vishnu
4 years ago
1 Mole =6.022*10^23N
So, in ½ or0.5 moles of H3O has 1/2*6.022*10^23*
N are there so ans is 1.5*10^21
Ans A
aswanth nayak
4 years ago
Dear Student,
 
 
1 Mole =6.022*10^23N
So, in ½ or0.5 moles of H3O has 1/2*6.022*10^23*
N are there so ans is 1.5*10^21
Correct Answer is option A